OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term 4x2 + ___x + 49

OpenStudy (vishweshshrimali5):

You must use this formula (a+b)^2 = a^2 + b^2 + 2ab Now compare ur eqn. with RHS of the above eq. Now you can find the value of missing term

OpenStudy (anonymous):

it would be 2*2*7=28

OpenStudy (vishweshshrimali5):

Vshal u r right dude

OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. x2 − 16x + ____ ?????

OpenStudy (vishweshshrimali5):

Now use this form (a-b)^2 = a^2 + b^2 - 2ab